At the beginning of 2019, Robotics Inc. acquired a manufacturing facility for $13.5 million. $10.5 million of the purchase price was allocated to the building. Depreciation for 2019 and 2020 was calculated using the straight-line method, a 25-year useful life, and a $2.5 million residual value. In 2021, the estimates of useful life and residual value were changed to 20 total years and $650,000, respectively. What is depreciation on the building for 2021?  
Answer #1

Depreciation of 2019 & 2020 on Building = (cost - residual value) / life of the assets

= (10500000 - 2500000)/25 = $320000

Net book value at the beginning of 2021 = 10500000 - 320000*2 = $9860000

Depreciation on the building for 2021 = (9860000 - 650000)/ 18 = $511667